OpenStudy (richyw):

surface area element. I am trying to remember how to find the surface area element of one of the faces of this pyramid.

OpenStudy (richyw):



OpenStudy (richyw):



OpenStudy (richyw):

so I am looking for \(d\vec{a}\). I know that this plane can be written as \[\frac{x}{a}+\frac{y}{b}+\frac{z}{c}=1\]

OpenStudy (richyw):

so I know how to get a vector normal to this surface. just a little bit unclear what how to turn that into "\(d\vec{a}\)"

OpenStudy (anonymous):

surface area element can be calculated ( lets call you surface f) |(div f)|dxdy / |[(div f) dot (z direction)]| if you want it as a vector you multiply it by the normal.